# TACLR: A Scalable and Efficient Retrieval-based Method for Industrial Product Attribute Value Identification

### v2

@@ Abstract
Product Attribute Value Identification (PAVI) involves identifying attribute values from product profiles, a key task for improving product search, recommendation, and business analytics on e-commerce platforms. However, existing PAVI methods face critical challenges, such as inferring implicit values, handling out-of-distribution (OOD) values, and producing normalized outputs. To address these limitations, we introduce Taxonomy-Aware Contrastive Learning Retrieval (TACLR), the first retrieval-based method for PAVI. TACLR formulates PAVI as an information retrieval task by encoding product profiles and candidate values into embeddings and retrieving values based on their similarity to the item embedding. It leverages contrastive training with taxonomy-aware hard negative sampling and employs adaptive inference with dynamic thresholds. TACLR offers three key advantages: (1) it effectively handles implicit and OOD values while producing normalized outputs; (2) it scales to thousands of categories, tens of thousands of attributes, and millions of values; and (3) it supports efficient inference for high-load industrial deployment. Extensive experiments on proprietary and public datasets validate the effectiveness and efficiency of TACLR. Moreover, it has been successfully deployed in a real-world e-commerce platform, processing millions of product listings daily while supporting dynamic, large-scale attribute taxonomies.
@@
Yindu Su1,2, Huike Zou1, Lin Sun3, Ting Zhang2, Haiyang Yang1, Liyu Chen1, David Lo2, Qingheng Zhang1, Shuguang Han1, Jufeng Chen1, 1Alibaba Group, 2Singapore Management University, 3Hangzhou City University, yindusu@foxmail.com qingheng.zqh@alibaba-inc.com, shuguang.sh@alibaba-inc.com
@@ 1 Introduction
Figure 1: Illustration of the PAVE and PAVI tasks. While an additional normalization step can adapt PAVE methods for PAVI, these methods remain unable to identify implicit values, such as Apple.
@@
Product attribute values are critical components that support the function of e-commerce platforms. They provide essential structural information, aiding customers in making informed purchasing decisions and enabling product listing Chen et al. (2024), recommendation Truong et al. (2022); Sun et al. (2020), retrieval 10.1145/3308560.3316603; Huang et al. (2014), and question answering Kulkarni et al. (2019); Gao et al. (2019).
@@
However, seller-provided attribute values are often incomplete or even inaccurate. This undermines the effectiveness of applications that rely on this information. Consequently, the automatic identification of product attribute values has become a critical challenge. Researchers have explored the task of Product Attribute Value Extraction (PAVE), which extracts spans from product profiles using Named Entity Recognition (NER) Zheng et al. (2018) or Question Answering (QA) Wang et al. (2020) models. The upper part of Figure 1 illustrates an example of NER-based PAVE.
@@
Although these approaches effectively extract value spans, the outputs remain raw subsequences. Presenting attribute values in a standardized format is crucial for facilitating data aggregation in business analytics and enhancing the user experience by providing clear and consistent information. To produce standardized values, a normalization step Putthividhya and Hu (2011); Zhang et al. (2021) is required to map these spans to predefined formats, as shown in the lower part of Figure 1. However, implicit values, such as Apple, cannot be directly extracted and must instead be inferred from context or prior knowledge.
@@
Therefore, in this work, we focus on the task of Product Attribute Value Identification (PAVI) Shinzato et al. (2023), which aims to associate predefined attribute values from attribute taxonomy (illustrated in Figure 2) with products. The input to PAVI includes the product category and profile, where the profile comprises textual data, such as the title and description, and may optionally include visual information, such as images or videos. The output is a dictionary with predefined attributes as keys and the inferred attribute values as corresponding entries. In addition, PAVI requires determining when attribute values are missing. For instance, as shown in Figure 1, the value for Version is unavailable and is therefore assigned an empty result or null value.
@@
Beyond adapting extraction-based approaches, researchers have investigated classification-based Chen et al. (2022) and generation-based paradigms Sabeh et al. (2024b) for PAVI. Classification-based methods treat each value as a label and employ multi-label classification models to recognize values across multiple attributes. While straightforward, these methods are fundamentally limited by their inability to identify out-of-distribution (OOD) values not present in the training data, making them impractical for the dynamic and continuously evolving nature of e-commerce platforms. In contrast, generation-based methods reformulate PAVI as a sequence-to-sequence problem. Although these methods can handle implicit and OOD values, they face significant challenges, such as generating uncontrollable outputs and incurring substantial computational costs in high-load scenarios due to their reliance on Large Language Models (LLMs).
@@
In summary, existing approaches face distinct challenges, including difficulties identifying implicit values, generalizing to OOD values, producing normalized outputs, or ensuring scalability and efficiency for large-scale industrial applications. To address these limitations, we propose a novel retrieval-based method, Taxonomy-Aware Contrastive Learning Retrieval (TACLR). Our approach formulates PAVI as an information retrieval task: the product item serves as the query, and the attribute taxonomy acts as the corpus, enabling the efficient retrieval of relevant attribute values as matched documents.
We use a shared encoder to generate embeddings for the input product and candidate values from the attribute taxonomy. The cosine similarity between these embeddings is computed and normalized to produce a relevance score. Our method adopts a contrastive learning framework inspired by CLIP Radford et al. (2021). Rather than relying on in-batch negatives, we implement taxonomy-aware negative sampling, which selects hard negative values from the same category and attribute to generate a more challenging and precise contrastive signal. Additionally, learnable null values are explicitly incorporated as the ground truth for product-attribute pairs without associated values. During inference, we address the limitations of static thresholds by introducing dynamic thresholds derived from the relevance score of null values. This adaptive inference mechanism improves the model’s ability to generalize across a large-scale attribute taxonomy.
Our contributions are threefold: (1) We propose a novel retrieval-based paradigm for PAVI, introducing a scalable and efficient framework capable of handling implicit values, generalizing to OOD values, and producing normalized outputs. (2) We incorporate contrastive training into TACLR, using a taxonomy-aware negative sampling strategy to improve representation discrimination. Additionally, TACLR features an adaptive inference mechanism that dynamically balances precision and recall in large-scale industrial applications. (3) We validate the effectiveness of TACLR through extensive experiments on proprietary and public datasets. In addition, TACLR has been successfully deployed in a real-world industrial environment, processing millions of product listings and supporting thousands of categories and millions of attribute values.
@@ 2.1 Product Attribute Value Extraction
While NER- and QA-based paradigms have proven effective for PAVE, they struggle to identify implicit attribute values. Additionally, both paradigms rely on post-extraction normalization to standardize values, using either string-based methods Putthividhya and Hu (2011) or semantic-based techniques Zhang et al. (2021). Furthermore, QA-based methods require processing a single product multiple times to handle multiple target attributes, leading to inefficiencies in large-scale settings. These limitations underscore the need for novel paradigms integrating extraction and normalization while addressing implicit values.
@@ 2.2 Product Attribute Value Identification
Classification-Based PAVI. A straightforward approach is to frame PAVI as a multi-label classification problem over a finite set of values. Chen et al. (2022) trained a unified classification model that masks invalid labels based on the product category. However, a significant limitation of this classification-based paradigm is its inability to recognize OOD values not included in the training set. This limitation reduces its practicality in dynamic e-commerce environments, where new categories and values frequently emerge.
@@
Figure 3: Overview of the training and inference process of our retrieval-based PAVI method. The left section illustrates contrastive training with taxonomy-aware negative sampling, while the right section demonstrates adaptive inference with pre-computed value embeddings.
@@ 3 Taxonomy-Aware Contrastive Learning Retrieval
This section defines the PAVI task with an attribute taxonomy (§3.1) and presents our retrieval-based paradigm for PAVI (§3.2). We then detail the use of contrastive training with taxonomy-aware negative sampling (§3.3) and an adaptive inference mechanism with dynamic thresholds (§3.4). Figure 3 provides an overview of the approach.
@@ 3.1 PAVI Task Definition
PAVI is grounded in an attribute taxonomy that encompasses numerous product categories. For each category $c$, the taxonomy specifies a set of attributes $\mathcal{A}_{c}=\{a_{1},a_{2},\dots\}$ relevant to products in that category, and for each attribute $a\in\mathcal{A}_{c}$, it provides a predefined set of standard values $\mathcal{V}_{a}=\{v_{1},v_{2},\dots\}$. Figure 2 illustrates this structure.
For a given product item $i$, with its title $t$ and description $d$, the item is assigned to a specific category $c$ with associated attributes $\mathcal{A}_{c}$. The objective of the PAVI task is to identify a relevant set of values $\mathcal{V}_{a}^{+}\subseteq\mathcal{V}_{a}$ for each attribute $a\in\mathcal{A}_{c}$. The set $\mathcal{V}_{a}^{+}$ can take one of three forms: a singleton ($\{v\}$), multiple values ($\{v_{1},v_{2},\dots\}$), or an empty set ($\varnothing$) if no information about $a$ is available in the product profile. Notably, a standard value may not always appear explicitly as a text span in $t$ or $d$; it may be conveyed in other forms. When a value is not explicitly mentioned, such cases are referred to as implicit values.
@@ 3.2 Retrieval-Based PAVI
To achieve this, we use encoders to generate embeddings for both the item and its candidate values. The cosine similarity between the item embedding and each candidate value embedding is computed and normalized to the range $[0,1]$ to measure relevance. For each attribute, the candidate values are ranked based on their similarity scores, and the most relevant values are selected as the output set.
To effectively encode both the item and candidate values, we preprocess them as textual inputs and use a shared text encoder. For each item, we concatenate its title $t$ and description $d$ as the input text. Each candidate value $v$ associated with attribute $a$ under category $c$ is represented as a text prompt, such as “A [phone ($c$)] with [brand ($a$)] being [Apple ($v$)]”.  [^1] This framework can be extended to multimodal scenarios by replacing the text encoder with a multimodal encoder to incorporate features like images. We explore the impact of various value prompt templates in §5.3.
@@ 3.3 Contrastive Training
$\mathcal{L}_{a}=-\log\left(\frac{\exp(\frac{s(i,v_{a}^{+})}{\tau})}{\exp(\frac{s(i,v_{a}^{+})}{\tau})+\sum\limits_{v\in\mathcal{V}_{a}^{-}}\exp(\frac{s(i,v)}{\tau})}\right),$
@@
$\mathcal{L}_{i}=\sum_{a\in\mathcal{A}_{c}}\mathcal{L}_{a}.$
@@ 3.4 Adaptive Inference
$\mathcal{V}_{a}^{\text{pred}}=\{v\mid s(i,v)>T^{\prime}_{a}\}.$
@@
$v_{a}^{\text{pred}}=\begin{cases}\arg\max\limits_{v\in\mathcal{V}_{a}}s(i,v)&\text{if }\max\limits_{v\in\mathcal{V}_{a}}s(i,v)>T^{\prime}_{a}\\ \textit{null}&\text{otherwise}\end{cases}.$
The inference process is illustrated on the right side of Figure 3, demonstrating how candidate value embeddings from the attribute taxonomy are pre-computed and stored offline to enhance efficiency. During online inference, the item profile is encoded into an item embedding $I$, which is then compared against groups of candidate value embeddings for various attributes. In this example, the predictions for $a_{3}$ and $a_{5}$ are determined to be empty because the highest-scoring value for these attributes is the null value.
@@ 4 Experiment Settings
Table 2: Statistics of the attribute taxonomies and dataset splits from Ecom-PAVI and WDC-PAVE. “CA Pairs” refers to category-attribute pairs, “CAV Tuples” denotes category-attribute-value tuples, “PA Pairs” represents product-attribute pairs, and “Null Pairs” indicate product-attribute pairs with null values. “Excl.” refers to the test split excluding measurement attributes.
@@
Statistic Ecom-PAVI WDC-PAVE
@@ 4.1 Datasets
To evaluate PAVI under the settings described in §3.1, we benchmark our proposed method against baselines using both proprietary and public datasets with normalized values. [^2] Other popular benchmarks such as AE-110k Xu et al. (2019) and MAVE Yang et al. (2022) provide only unnormalized values as spans extracted from product profiles, making them unsuitable for our experiments. Table 2 presents statistics of the attribute taxonomies and the datasets.
Ecom-PAVI. This dataset, derived from a real-world e-commerce platform, is designed to evaluate the scalability and generalization of PAVI methods. The attribute taxonomy in the e-commerce platform comprises 8,803 product categories, 26,645 category-attribute pairs, and 6.3 million category-attribute-value tuples. For our experiments, we sampled 1 million products for training, 10,000 for validation, and 10,000 for testing, ensuring that the samples span different time periods to reflect real-world scenarios. To ensure data quality, annotators manually verified the assigned product categories, discarded incorrectly categorized products, and selected the corresponding attribute-value pairs from the taxonomy as the ground truth.
WDC-PAVE Brinkmann et al. (2024). This dataset consists of products distributed across 5 categories. The training set includes 1,066 products and 8,832 product-attribute pairs, of which 3,973 have null values. The test set contains 354 products and 2,937 product-attribute pairs, with 1,330 null pairs. We conduct two evaluations: the first on the original test set, which includes all attributes, and the second on a test split that excludes measurement attributes to focus on tasks not requiring complex reasoning for unit conversion.

### v4

@@ Abstract
Product Attribute Value Identification (PAVI) involves identifying attribute values from product profiles, a key task for improving product search, recommendation, and business analytics on e-commerce platforms. However, existing PAVI methods face critical challenges, such as inferring implicit values, handling out-of-distribution (OOD) values, and producing normalized outputs. To address these limitations, we introduce Taxonomy-Aware Contrastive Learning Retrieval (TACLR), the first retrieval-based method for PAVI. TACLR formulates PAVI as an information retrieval task by encoding product profiles and candidate values into embeddings and retrieving values based on their similarity. It leverages contrastive training with taxonomy-aware hard negative sampling and employs adaptive inference with dynamic thresholds. TACLR offers three key advantages: (1) it effectively handles implicit and OOD values while producing normalized outputs; (2) it scales to thousands of categories, tens of thousands of attributes, and millions of values; and (3) it supports efficient inference for high-load industrial deployment. Extensive experiments on proprietary and public datasets validate the effectiveness and efficiency of TACLR. Further, it has been successfully deployed on the real-world e-commerce platform Xianyu, processing millions of product listings daily with frequently updated, large-scale attribute taxonomies. We release the code to facilitate reproducibility and future research at https://github.com/SuYindu/TACLR.
@@
Yindu Su1,2, Huike Zou1, Lin Sun3, Ting Zhang2, Haiyang Yang1, Liyu Chen1, David Lo2, Qingheng Zhang1, Shuguang Han1, Jufeng Chen1, 1Xianyu of Alibaba  2Singapore Management University  3Hangzhou City University yindusu@foxmail.com, tingzhang.2019@phdcs.smu.edu.sg qingheng.zqh@alibaba-inc.com, shuguang.sh@alibaba-inc.com
@@ 1 Introduction
Figure 1: Illustration of the PAVE and PAVI tasks. Unlike PAVE, which extracts raw value spans from product profiles, PAVI requires outputs to be normalized and supports both the identification of implicit values and the assignment of null to unavailable attribute values.
@@
Product attribute values are key components that support the operation of e-commerce platforms. They provide essential structural information, aiding customers in making informed purchasing decisions and enabling product listing Chen et al. (2024), recommendation Truong et al. (2022); Sun et al. (2020), retrieval Magnani et al. (2019); Huang et al. (2014), and question answering Kulkarni et al. (2019); Gao et al. (2019).
@@
However, seller-provided attribute values are often incomplete or even inaccurate—an issue that is particularly severe on second-hand e-commerce platforms such as Xianyu  [^1] https://www.goofish.com. This undermines the effectiveness of downstream applications, making the automatic identification of product attribute values a fundamental requirement. Researchers have explored Product Attribute Value Extraction (PAVE), which involves extracting spans from product profiles using Named Entity Recognition (NER) Zheng et al. (2018) or Question Answering (QA) Wang et al. (2020) models. The upper part of Figure 1 illustrates an example of NER-based PAVE.
@@
Although these approaches effectively extract value spans, the outputs remain raw text subsequences. Presenting attribute values in a standardized format is crucial for facilitating data aggregation in business analytics and enhancing the user experience by providing clear and consistent information. To produce standardized values, a normalization step Putthividhya and Hu (2011); Zhang et al. (2021) is required to map these spans to predefined formats, as shown in the lower part of Figure 1. However, implicit values, such as Apple, cannot be directly extracted and must instead be inferred from context or prior knowledge.
@@
Therefore, in this work, we focus on Product Attribute Value Identification (PAVI) Shinzato et al. (2023), which aims to associate predefined attribute values from attribute taxonomy (illustrated in Figure 2) with products. The input for PAVI includes the product category and profile, where the profile includes textual data, such as the title and description, and may optionally incorporate visual information, such as images or videos. The output is a dictionary with predefined attributes as keys and their inferred values as corresponding entries. In addition, PAVI requires determining when attribute values are missing. For instance, as shown in Figure 1, the value for Version is unavailable and is therefore assigned an empty result or null value.
@@
Beyond adapting PAVE approaches, researchers have investigated classification-based Chen et al. (2022) and generation-based paradigms Sabeh et al. (2024b) for PAVI. Classification-based methods treat each value as a class; while this approach is straightforward, it is fundamentally limited by the inability to identify out-of-distribution (OOD) values not present in the training data, making such methods impractical for the continuously evolving nature of e-commerce platforms. In contrast, generation-based methods reformulate PAVI as a sequence-to-sequence task. Although these methods can handle implicit and OOD values, they face significant challenges, such as generating uncontrollable outputs and incurring substantial computational costs in high-load scenarios due to their reliance on Large Language Models (LLMs). In summary, existing approaches face distinct challenges, including difficulties in identifying implicit values, generalizing to OOD values, producing normalized outputs, and ensuring scalability and efficiency for large-scale industrial applications.
@@
To address these limitations, we propose a novel retrieval-based method, Taxonomy-Aware Contrastive Learning Retrieval (TACLR). Our approach formulates PAVI as an information retrieval task: the product item serves as the query, and the attribute taxonomy acts as the corpus, enabling efficient retrieval of relevant attribute values as matched documents. We use a shared encoder to generate embeddings for both the input product and candidate values from the attribute taxonomy. Our method adopts a contrastive learning framework inspired by CLIP Radford et al. (2021). Rather than relying on in-batch negatives, we implement taxonomy-aware negative sampling, which selects hard negative values from the same category and attribute to generate a more challenging and precise contrastive signal. Additionally, learnable null values are explicitly incorporated as the ground truth for product-attribute pairs without associated values. During inference, we address the limitations of static thresholds by introducing dynamic thresholds derived from the relevance score of null values. This adaptive inference mechanism improves the model’s ability to generalize across a large-scale attribute taxonomy.
Our contributions are threefold: (1) We propose a novel retrieval-based paradigm for PAVI, introducing a scalable and efficient framework capable of handling implicit values, generalizing to OOD values, and producing normalized outputs. (2) We incorporate contrastive training into TACLR, using a taxonomy-aware negative sampling strategy to improve representation discrimination, and introduce an adaptive inference mechanism that dynamically balances precision and recall in large-scale industrial applications. (3) We validate the effectiveness of TACLR through extensive experiments on proprietary and public datasets, and demonstrate its successful deployment in a real-world industrial environment, processing millions of product listings across thousands of categories and millions of attribute values.
@@ 2.1 Product Attribute Value Extraction
While NER- and QA-based paradigms have proven effective for PAVE, they struggle to identify implicit attribute values. Additionally, both paradigms rely on post-extraction normalization to standardize values, using either lexical Putthividhya and Hu (2011) or semantic methods Zhang et al. (2021). Furthermore, extraction-based models require token-level annotations (e.g., BIO tags) for training and evaluation. Producing such annotations is significantly more resource-intensive than generating the value-level annotations used by TACLR, further limiting the scalability of these extraction-based methods.
@@ 2.2 Product Attribute Value Identification
Classification-Based PAVI. A straightforward approach is to frame PAVI as a multi-label classification problem over a finite set of values. Chen et al. (2022) trained a unified classification model that masks invalid labels based on the product category. However, a significant limitation of this classification-based paradigm is its inability to recognize OOD values not included in the training set. Consequently, its practicality is limited in dynamic e-commerce environments, where new categories and values frequently emerge.
@@
Figure 3: Overview of the training and inference process of TACLR, our retrieval-based method for the PAVI task. The left section illustrates contrastive training with taxonomy-aware negative sampling, while the right section demonstrates adaptive inference with pre-computed value embeddings.
@@ 3 Taxonomy-Aware Contrastive Learning Retrieval
This section defines the PAVI task with an attribute taxonomy (§3.1) and presents our retrieval-based paradigm for PAVI (§3.2). We then detail the use of contrastive training with taxonomy-aware negative sampling (§3.3) and an adaptive inference mechanism with dynamic thresholds (§3.4). Figure 3 provides an overview of our approach TACLR.
@@ 3.1 PAVI Task Definition
PAVI is grounded in an attribute taxonomy that encompasses numerous product categories. For each category $c$, the taxonomy specifies a set of attributes $\mathcal{A}_{c}=\{a_{1},a_{2},\dots\}$ relevant to products in that category. For each attribute $a\in\mathcal{A}_{c}$, it provides a predefined set of standard values $\mathcal{V}_{a}=\{v_{1},v_{2},\dots\}$. Figure 2 illustrates this structure [^2] Our approach focuses on attribute value identification, leveraging an existing attribute taxonomy as input rather than constructing or updating the taxonomy itself. In the Xianyu platform, a dedicated team and supporting system are responsible for maintaining the taxonomy (i.e., “attribute mining”)..
For a given product item $i$, with its title $t$ and description $d$, the item is assigned to a category $c$ with associated attributes $\mathcal{A}_{c}$. The objective of the PAVI task is to identify a relevant set of values $\mathcal{V}_{a}^{+}\subseteq\mathcal{V}_{a}$ for each attribute $a\in\mathcal{A}_{c}$. The set $\mathcal{V}_{a}^{+}$ can take one of three forms: a singleton ($\{v\}$), multiple values ($\{v_{1},v_{2},\dots\}$), or an empty set ($\varnothing$) if no information about $a$ is available in the product profile. Notably, a standard value may not always appear verbatim as a text span in $t$ or $d$; it may instead be present in a paraphrased or synonymous form, which we refer to as an unnormalized value (e.g., the standard value iPhone 12 Pro Max expressed as iphone12pm in Figure 1). In other cases, a value may not be explicitly mentioned in the product profile but can be inferred from the context; these are referred to as implicit values (e.g., Apple in Figure 1).
@@ 3.2 Retrieval-Based PAVI
Encoding of items and values. We preprocess both item profiles and candidate values as textual inputs, utilizing a shared text encoder. For each item, we concatenate its title ($t$) and description ($d$) into a single input sentence formatted as: title: {title} description: {description}. Each candidate value is represented as a context-rich prompt, structured as: A {category} with {attribute} being {value}, e.g., A phone with brand being Apple. We explore the impact of various prompt templates in §5.3  [^3] This framework can be extended to multimodal scenarios by replacing the text encoder with a multimodal encoder to incorporate features such as images..
Inference pipeline. During the deployment process, all value embeddings within the attribute taxonomy are pre-computed offline and indexed using the Faiss library [^4] https://github.com/facebookresearch/faiss. During online inference, each item is encoded into an embedding, which is then compared against groups of indexed candidate value embeddings for various attributes. For each attribute $a\in\mathcal{A}_{c}$, the top-$k$ most similar values are retrieved whose similarity scores exceed an adaptive threshold (§3.4).
@@ 3.3 Contrastive Training
$\mathcal{L}_{a}=-\log\left(\frac{\exp(\frac{s(i,v_{a}^{+})}{\tau})}{\exp(\frac{s(i,v_{a}^{+})}{\tau})+\sum\limits_{v\in\mathcal{V}_{a}^{-}}\exp(\frac{s(i,v)}{\tau})}\right)$
@@
$\mathcal{L}_{i}=\sum_{a\in\mathcal{A}_{c}}\mathcal{L}_{a}$
@@ 3.4 Adaptive Inference
$\mathcal{V}_{a}^{\text{pred}}=\{v\mid s(i,v)>T^{\prime}_{a}\}$
@@
$v_{a}^{\text{pred}}=\begin{cases}\arg\max\limits_{v\in\mathcal{V}_{a}}s(i,v)&\text{if }\max\limits_{v\in\mathcal{V}_{a}}s(i,v)>T^{\prime}_{a}\\ \text{null}&\text{otherwise}\end{cases}$
Equivalently, we can include the null value as an explicit candidate and retrieve the top-1 value as:
$v_{a}^{\text{pred}}=\arg\max\limits_{v\in\mathcal{V}_{a}\cup\{v_{0}^{a}\}}s(i,v)$
In this case, selecting the $v_{0}^{a}$ corresponds to the scenario where none of the specific candidate values surpass the dynamic threshold.
The inference process is illustrated on the right side of Figure 3. In this example, the model retrieves $v_{3}^{1}$, $v_{1}^{2}$, and $v_{2}^{4}$ as the predicted values for attributes $a_{1}$, $a_{2}$, and $a_{4}$, respectively, while those for $a_{3}$ and $a_{5}$ are determined to be empty as their top-scoring values are null.
@@ 4 Experiment Settings
Table 2: Statistics of the attribute taxonomies and dataset splits from Xianyu-PAVI and WDC-PAVE. “CA Pairs” refers to category-attribute pairs, “CAV Tuples” denotes category-attribute-value tuples, “PA Pairs” represents product-attribute pairs, and “Null Pairs” indicates product-attribute pairs with null values. “Excl.” refers to the test split excluding measurement attributes.
@@
Statistic Xianyu-PAVI WDC-PAVE
@@ 4.1 Datasets
To evaluate PAVI under the settings described in §3.1, we compare our proposed method against baselines on both proprietary and public datasets with normalized values [^5] Other popular benchmarks, such as AE-110k Xu et al. (2019) and MAVE Yang et al. (2022), provide only unnormalized values as spans extracted from product profiles, making them unsuitable for our experiments.. Table 2 presents statistics of the attribute taxonomies and datasets.
Xianyu-PAVI. This dataset, derived from the e-commerce platform Xianyu, is constructed to evaluate the scalability and generalization of PAVI methods. The platform’s attribute taxonomy comprises 8,803 product categories, 26,645 category-attribute pairs, and 6.3 million category-attribute-value tuples. On average, each category is associated with 3 attributes, each attribute has 237 possible values, and there are 716 candidate values per category.
For our experiments, we randomly sampled 1 million product items for training, 10,000 for validation, and 10,000 for testing. Each item is annotated with a category label and multiple attribute-value pair labels. Category labels were generated through a multi-step process involving automated classification, seller feedback, and annotator review, during which misclassified samples were discarded. Attribute-value pair labels were obtained through a multi-stage manual annotation process: a pool of annotators conducted the initial labeling, followed by quality checks and a second round of review, with items reassigned to different annotators after shuffling.
WDC-PAVE Brinkmann et al. (2024). This dataset contains 1,066 training and 354 test product items across 5 categories, with 8,832 and 2,937 product-attribute pairs (3,973 and 1,330 nulls), respectively. On average, each category is associated with 7.4 attributes, each attribute has 62 values, and there are 459 attribute-value pairs per category. We conduct two evaluations: the first on the original test set, and the second on a test split that excludes measurement attributes, which require complex reasoning for unit conversion.